1)The difference of three times x and y is 11. Please translate into an equation.
2)y less than 3 times x is 9. Again, please translate into an equation.

Found 2 solutions by stanbon, Jenny_has_lost_her_mind:
Answer by stanbon(75887)

1)The difference of three times x and y is 11.
3x - y = 11
----------------------
Please translate into an equation.
2)y less than 3 times x is 9
3x - y = 9
==========================
Cheers,
Stan H.

Answer by Jenny_has_lost_her_mind(1)

1) 3x-y=11
2) 3x-y=9